Solve the equation a = m - n for the variable n.
A. n = a + m
B. n = m - a
C. n = a - m
D. n = - m - a

Answer by dkppathak(439)

Solve the equation a = m - n for the variable n.
A. n = a + m
B. n = m - a
C. n = a - m
D. n = - m - a
solution
a=m-n where n is variable , we have to find the value of n . keeping n positive shift to other side and a to other side n negative change to positive and a change to negative
so equation will be
n=m-a
option B is right answer